OpenStudy (anonymous):

6:11 is in propotion with 10:22. True or False

OpenStudy (anonymous):

Set them up as a proportion and find the cross products: 6/11 = 10/22 6 x 22 = 10 x 11? if so, it is true; if not, it is false

OpenStudy (anonymous):